# Unraveling Zeroth-Order Optimization through the Lens of Low-Dimensional Structured Perturbations

### v1

@@ B.3 Technical Lemmas for Gaussian Smoothed Loss
The expected norm of structured zeroth-order gradient is bounded by
@@
under Lemma B.4 and Assumption 3.4. Therefore, we have
@@
where we use Lemma B.3 and Lemma B.5. Lastly, taking the expectations with respect to all randomness, we get
@@
$\leq\frac{15\mu^{2}L^{2}s^{3}}{2}+4(s+2)\left(\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+\frac{\sigma^{2}}{B}\right)$
@@
B.4 Proof of Theorem 3.6
By the $L$-smooth loss property, we have
$\mathcal{L}(\bm{\theta}_{t+1}+\mu\mathbf{u})$ $\leq\mathcal{L}(\bm{\theta}_{t}+\mu\mathbf{u})+\langle\nabla\mathcal{L}(\bm{\theta}_{t}+\mu\mathbf{u}),\bm{\theta}_{t+1}-\bm{\theta}_{t}\rangle+\frac{L}{2}\|\bm{\theta}_{t+1}-\bm{\theta}_{t}\|^{2}$
Taking the expectation for both sides yields that
$\underbrace{\mathbb{E}_{\mathbf{u}}\big{[}\mathcal{L}(\bm{\theta}_{t+1}+\mu\mathbf{u})\big{]}}_{\mathcal{L}_{\mu}(\bm{\theta}_{t+1})}$ $\leq\mathbb{E}_{\mathbf{u}}\Big{[}\mathcal{L}(\bm{\theta}_{t}+\mu\mathbf{u})+\langle\nabla\mathcal{L}(\bm{\theta}_{t}+\mu\mathbf{u}),\bm{\theta}_{t+1}-\bm{\theta}_{t}\rangle+\frac{L}{2}\|\bm{\theta}_{t+1}-\bm{\theta}_{t}\|^{2}\Big{]}$
$=\mathcal{L}_{\mu}(\bm{\theta}_{t})-\alpha\langle\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t}),\widehat{\mathbf{g}}_{t}\rangle+\frac{\alpha^{2}L}{2}\|\widehat{\mathbf{g}}_{t}\|^{2}$
Again, by expectation with respect to $\mathbf{u}_{t}$, we have
$\mathcal{L}_{\mu}(\bm{\theta}_{t+1})$ $\leq\mathcal{L}_{\mu}(\bm{\theta}_{t})-\alpha\big{\langle}\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t}),\nabla\mathcal{L}_{\mu,\mathbf{M}_{t}}(\bm{\theta}_{t})\big{\rangle}+\frac{\alpha^{2}L}{2}\mathbb{E}_{\mathbf{u}_{t}}\big{[}\|\widehat{\mathbf{g}}_{t}\|^{2}\big{]}$
$\leq\mathcal{L}_{\mu}(\bm{\theta}_{t})-\alpha\Bigg{(}\|\nabla\mathcal{L}_{\mu}(\bm{\theta})\|^{2}-\|\nabla\mathcal{L}_{\mu}(\bm{\theta})\|\mu L\sqrt{2(d+s)}\Bigg{)}+\frac{\alpha^{2}L}{2}\mathbb{E}_{\mathbf{u}_{t}}\big{[}\|\widehat{\mathbf{g}}_{t}\|^{2}\big{]}$
$\leq\mathcal{L}_{\mu}(\bm{\theta}_{t})-\alpha\|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})\|^{2}+\frac{\alpha}{2}\|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})\|^{2}+\alpha\mu^{2}L^{2}(d+s)+\frac{\alpha^{2}L}{2}\mathbb{E}_{\mathbf{u}_{t}}\big{[}\|\widehat{\mathbf{g}}_{t}\|^{2}\big{]}$
By Lemma B.7, we have
$\mathcal{L}_{\mu}(\bm{\theta}_{t+1})$ $\leq\mathcal{L}_{\mu}(\bm{\theta}_{t})-\frac{\alpha}{2}\|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})\|^{2}+\frac{\alpha\mu^{2}L^{2}R}{2}+\frac{\alpha^{2}L}{2}\Bigg{(}\frac{15\mu^{2}L^{2}s^{3}}{2}+2(s+2)\left(\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+\frac{\sigma^{2}}{B}\right)\Bigg{)}$
Rearranging the inequality, we obtain
$\frac{\alpha}{4}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ $\leq\frac{\alpha\mu^{2}L^{2}(d+3)^{3}}{8}+\frac{\alpha}{2}\|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})\|^{2}$
$\leq\frac{\alpha\mu^{2}L^{2}(d+3)^{3}}{8}+\mathcal{L}_{\mu}(\bm{\theta}_{t})-\mathcal{L}_{\mu}(\bm{\theta}_{t+1})+\alpha\mu^{2}L^{2}(d+s)+\frac{15\alpha^{2}\mu^{2}L^{3}s^{3}}{4}$
$\quad+\alpha^{2}L(s+2)\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+\frac{\alpha^{2}L(s+2)\sigma^{2}}{B}$
We assume the following stepsize condition
$\alpha$ $\leq\frac{1}{8L(s+2)}.$
Finally, we have
$\frac{\alpha}{8}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ $\leq\frac{\alpha\mu^{2}L^{2}(d+3)^{3}}{8}+\mathcal{L}_{\mu}(\bm{\theta}_{t})-\mathcal{L}_{\mu}(\bm{\theta}_{t+1})+\alpha\mu^{2}L^{2}(d+s)+\frac{15\alpha^{2}\mu^{2}L^{3}s^{3}}{4}+\frac{\alpha^{2}L(s+2)\sigma^{2}}{B}$
Hence, we obtain
$\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ $\leq\mu^{2}L^{2}(d+3)^{3}+\frac{8\big{(}\mathcal{L}_{\mu}(\bm{\theta}_{t})-\mathcal{L}_{\mu}(\bm{\theta}_{t+1})\big{)}}{\alpha}+8\mu^{2}L^{2}(d+s)+30\alpha\mu^{2}L^{3}s^{3}+\frac{8L(s+2)\sigma^{2}}{B}$
We let $\Delta=\mathcal{L}_{\mu}(\bm{\theta}_{1})-\mathcal{L}_{\mu}(\bm{\theta}^{*})$. Under the smoothing parameter condition $\mu=\mathcal{O}(\frac{1}{L\sqrt{sdT}})$, we finally arrive at
$\mathbb{E}\Big{[}\frac{1}{T}\sum\limits_{t=1}^{T}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}\Big{]}$ $\leq\mu^{2}L^{2}(d+3)^{3}+\frac{8\Delta}{\alpha T}+8\mu^{2}L^{2}(d+s)+30\alpha\mu^{2}L^{3}s^{3}+\frac{8\alpha L(s+2)\sigma^{2}}{B}$
$\leq\Big{[}\mu^{2}L^{2}d^{3}+\frac{\Delta}{\alpha T}+\mu^{2}L^{2}s^{2}\Big{]}\mathcal{O}\Big{(}\frac{1}{T}\Big{)}+\mathcal{O}\Big{(}\frac{\sigma^{2}}{B}\Big{)}$
$\leq\mathcal{O}\Big{(}\frac{d^{2}}{sT}+\frac{s^{2}}{dT}+\frac{\Delta}{\alpha T}+\frac{\sigma^{2}}{B}\Big{)}$
B.5 Proof of Theorem 3.10
In this section, we provide the convergence analysis with additional assumption on the intrinsic dimension of the loss function.
@@
Let $\mathcal{L}_{\mu}(\bm{\theta})=\mathbb{E}\big{[}\mathcal{L}(\bm{\theta}+\mu\mathbf{u})\big{]}$ for some positive $\mu>0$. Then, with locally low intrinsic dimension, the following holds.
$\|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})-\nabla\mathcal{L}(\bm{\theta}_{t})\|$ $\leq\frac{\sqrt{3}\mu Lr\sqrt{d}}{2}+\frac{\mu L(d+4)\sqrt{d}}{2}\exp(-\frac{c_{1}}{2}d)$
$\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ $\leq 3\mu^{2}L^{2}r^{2}d+\mu^{2}L^{2}d(d+4)^{2}\exp(-c_{1}d)+2|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})\|^{2}$
@@
Let $\mathcal{E}$ be the event such that $\|\bm{\theta}_{t}+\mu\mathbf{u}-\bm{\theta}_{t}\|\leq 2\alpha sG+2\mu\sqrt{d}$. On the event $\mathcal{E}$, by the low intrinsic dimension condition, there exists some matrix $\mathbf{H}(\bm{\theta}_{t})$ such that $\nabla^{2}\mathcal{L}(\bm{\theta}_{t}+\mu\mathbf{u})\preceq\mathbf{H}(\bm{\theta}_{t})$ with $\mathrm{intdim}(\mathbf{H}(\bm{\theta}_{t}))\leq r$. Thus, we have the following inequality
@@
Likewise, on the event $\mathcal{E}^{c}$, we have
@@
$\mathbb{E}\big{[}\|\mathbf{u}\|^{2}\big{]}$ $\leq d$
Therefore, we could obtain
@@
Hence, we further obtain
$\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ $\leq 2\|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})-\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+2\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$
@@
$\qquad+\frac{\alpha^{2}}{2}\Big{\langle}\mathbb{E}\Big{[}\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}\!\cdot\!\mathbb{I}(\mathcal{E}^{c})\Big{]},L\!\cdot\!\mathbf{I}_{d}-\mathbf{H}(\bm{\theta}_{t})\Big{\rangle},$
@@
Using Hanson-Wright inequality (Hanson & Wright, 1971; Wright, 1973), it is easy to check that $\mathbb{P}[\mathcal{E}^{c}]\leq\exp(-s/2)$ holds. Combining all the items, we finally have
@@
$\mathbb{E}\big{[}\widehat{\mathbf{g}}_{t}^{\top}\mathbf{H}(\bm{\theta}_{t})\widehat{\mathbf{g}}_{t}\big{]}$ $\leq 13\mu^{2}L^{3}s^{2}\rho+9L\rho\Big{(}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+\frac{\sigma^{2}}{B}\Big{)}.$
@@
Also, by the definition of effective overlap in Definition 3.9, we have
@@
$\leq 13\mu^{2}L^{3}s^{2}\rho_{t}+9L\rho_{\text{max}}\Big{(}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+\frac{\sigma^{2}}{B}\Big{)}$
Proof of Theorem 3.10
@@
We have
$\mathcal{L}_{\mu}(\bm{\theta}_{t+1})$ $\leq\mathcal{L}_{\mu}(\bm{\theta}_{t})-\frac{\alpha}{2}\|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})\|^{2}+\alpha\mu^{2}L^{2}(d+s)+\frac{\alpha^{2}}{2}\Big{(}13\mu^{2}L^{3}s^{2}\rho_{t}+9L\rho_{\text{max}}\Big{(}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+\frac{\sigma^{2}}{B}\Big{)}\Big{)}$
@@
By Lemma B.8, we have
@@
$\leq\frac{3\alpha\mu^{2}L^{2}r^{2}d}{16}+\mathcal{L}_{\mu}(\bm{\theta}_{t})-\mathcal{L}_{\mu}(\bm{\theta}_{t+1})+\alpha\mu^{2}L^{2}(d+s)+\frac{13\alpha^{2}\mu^{2}L^{3}s^{2}\rho_{t}}{2}+\frac{9\alpha^{2}L\rho_{\text{max}}}{2}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$
$\qquad+\frac{9\alpha^{2}L\rho_{\text{max}}\sigma^{2}}{2B}+37\alpha^{2}\mu^{2}L^{3}s^{3}\exp\big{(}-s/2\big{)}+42\alpha^{2}Ls\exp\big{(}-s/2\big{)}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$
@@
By rearranging inequality, we have
$\Big{(}\frac{\alpha}{4}-\frac{9\alpha^{2}L\rho_{\text{max}}}{2}-42\alpha^{2}Ls\exp(-s/2)\Big{)}\sum\limits_{t=1}^{T}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ $\leq\frac{3\alpha\mu^{2}L^{2}r^{2}d}{16}+\mathcal{L}_{\mu}(\bm{\theta}_{t})-\mathcal{L}_{\mu}(\bm{\theta}_{t+1})+\alpha\mu^{2}L^{2}(d+s)$
$\qquad+\frac{13\alpha^{2}L^{3}s^{2}\rho_{t}}{2}+\frac{9\alpha^{2}L\rho_{\text{max}}\sigma^{2}}{2B}+37\alpha^{2}\mu^{2}L^{3}s^{3}\exp(-s/2)$
@@
$\alpha\leq\frac{1}{36L\rho_{\text{max}}+336Ls\exp(-s/2)},$
$\frac{\alpha}{8}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ $\leq\frac{3\alpha\mu^{2}L^{2}r^{2}d}{16}+\mathcal{L}_{\mu}(\bm{\theta}_{t})-\mathcal{L}_{\mu}(\bm{\theta}_{t+1})+\alpha\mu^{2}L^{2}(d+s)+\frac{13\alpha^{2}\mu^{2}L^{3}s^{2}\rho_{t}}{2}+\frac{9\alpha^{2}L\rho_{\text{max}}\sigma^{2}}{2B}$
@@
$\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ $\leq\frac{3\mu^{2}L^{2}r^{2}d}{2}+\frac{8(\mathcal{L}_{\mu}(\bm{\theta}_{t})-\mathcal{L}_{\mu}(\bm{\theta}_{t+1}))}{\alpha}+8\mu^{2}L^{2}(d+s)+52\alpha\mu^{2}L^{3}s^{2}\rho_{t}+\frac{36\alpha L\rho_{\text{max}}\sigma^{2}}{B}$

### v2

@@ B.4 Technical Lemmas for Gaussian Smoothed Loss
Suppose that $\mathrm{srank}(\mathbf{M}_{t})\leq s$. Then, the expected norm of structured zeroth-order gradient satisfies
@@
under Lemma B.4 and $\mathrm{srank}(\mathbf{M}_{t})\leq s$. Hence, it follows that
@@
where we use Lemma B.3 and Lemma B.5. Lastly, taking the expectations with respect to all randomness, we obtain
@@
$\leq\frac{15\mu^{2}L^{2}s^{3}}{2}+4(s+2)\left(\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+\frac{\sigma^{2}}{B}\right),$
@@
From now, we provide some auxiliary lemmas under $\mathrm{srank}(\mathbf{M}_{t})\leq s$ and Assumption 3.1.
@@
The smoothed loss $\mathcal{L}_{\mu}$ under Assumption 3.1 satisfies
$\|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})-\nabla\mathcal{L}(\bm{\theta}_{t})\|$ $\leq\frac{\sqrt{3}\mu Lr\sqrt{d}}{2}+\frac{\mu L(d+4)\sqrt{d}}{2}\exp(-\frac{c_{1}}{2}d),$
$\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ $\leq 3\mu^{2}L^{2}r^{2}d+\mu^{2}L^{2}d(d+4)^{2}\exp(-c_{1}d)+2\|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})\|^{2}.$
@@
Let $\mathcal{E}$ be the event such that $\|\bm{\theta}_{t}+\mu\mathbf{u}-\bm{\theta}_{t}\|\leq 2\alpha sG+2\mu\sqrt{d}$. On the event $\mathcal{E}$, by Assumption 3.1, there exists some matrix $\mathbf{H}(\bm{\theta}_{t})$ such that $\nabla^{2}\mathcal{L}(\bm{\theta}_{t}+\mu\mathbf{u})\preceq\mathbf{H}(\bm{\theta}_{t})$ with $\mathrm{intdim}(\mathbf{H}(\bm{\theta}_{t}))\leq r$. Thus, we have the following inequality
@@
Likewise, on the event $\mathcal{E}^{c}$, it should hold that
@@
$\mathbb{E}\big{[}\|\mathbf{u}\|^{4}\big{]}$ $\leq(d+4)^{2}$
Therefore, it follows that
@@
Hence, we finally obtain
$\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ $\leq 2\|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})-\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+2\|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})\|^{2}$
@@
$\qquad\qquad+\frac{\alpha^{2}}{2}\Big{\langle}\mathbb{E}\Big{[}\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}\!\cdot\!\mathbb{I}(\mathcal{E}^{c})\Big{]},L\!\cdot\!\mathbf{I}_{d}-\mathbf{H}(\bm{\theta}_{t})\Big{\rangle},$
and the last term can be bounded by
@@
Using Hanson-Wright inequality (Hanson and Wright, 1971; Wright, 1973), it is easy to check that $\mathbb{P}[\mathcal{E}^{c}]\leq\exp(-s/2)$ holds. Combining all the items, we finally have
@@
$\mathbb{E}\big{[}\widehat{\mathbf{g}}_{t}^{\top}\mathbf{H}(\bm{\theta}_{t})\widehat{\mathbf{g}}_{t}\big{]}$ $\leq 13\mu^{2}L^{3}s^{2}\rho_{t}+9L\rho_{\mathrm{max}}\Big{(}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+\frac{\sigma^{2}}{B}\Big{)}.$
@@
Also, by the definition of subspace alignment in Definition 3.2, we have
@@
$\leq 13\mu^{2}L^{3}s^{2}\rho_{t}+9L\rho_{\mathrm{max}}\Big{(}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+\frac{\sigma^{2}}{B}\Big{)}$
B.5 Proof of Theorem 3.3
@@
By Lemma B.9 and Lemma B.10, it follows that
$\mathcal{L}_{\mu}(\bm{\theta}_{t+1})$ $\leq\mathcal{L}_{\mu}(\bm{\theta}_{t})-\frac{\alpha}{2}\|\nabla\mathcal{L}_{\mu}(\bm{\theta}_{t})\|^{2}+\alpha\mu^{2}L^{2}(d+s)+\frac{\alpha^{2}}{2}\Big{(}13\mu^{2}L^{3}s^{2}\rho_{t}+9L\rho_{\mathrm{max}}\Big{(}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+\frac{\sigma^{2}}{B}\Big{)}\Big{)}$
@@
By Lemma B.8, the gradient norm w.r.t. to the smoothed loss can be bounded by the gradient norm w.r.t. the original loss $\mathcal{L}$ as
@@
$\leq\frac{3\alpha\mu^{2}L^{2}r^{2}d}{16}+\mathcal{L}_{\mu}(\bm{\theta}_{t})-\mathcal{L}_{\mu}(\bm{\theta}_{t+1})+\alpha\mu^{2}L^{2}(d+s)+\frac{13\alpha^{2}\mu^{2}L^{3}s^{2}\rho_{t}}{2}+\frac{9\alpha^{2}L\rho_{\mathrm{max}}}{2}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$
$\qquad+\frac{9\alpha^{2}L\rho_{\mathrm{max}}\sigma^{2}}{2B}+37\alpha^{2}\mu^{2}L^{3}s^{3}\exp\big{(}-s/2\big{)}+42\alpha^{2}Ls\exp\big{(}-s/2\big{)}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$
@@
By rearranging inequality, it should hold that
$\Big{(}\frac{\alpha}{4}-\frac{9\alpha^{2}L\rho_{\mathrm{max}}}{2}-42\alpha^{2}Ls\exp(-s/2)\Big{)}$ $\sum\limits_{t=1}^{T}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}\leq$
$\frac{3\alpha\mu^{2}L^{2}r^{2}d}{16}+\mathcal{L}_{\mu}(\bm{\theta}_{t})-\mathcal{L}_{\mu}(\bm{\theta}_{t+1})+\alpha\mu^{2}L^{2}(d+s)$
$\qquad+\frac{13\alpha^{2}L^{3}s^{2}\rho_{t}}{2}+\frac{9\alpha^{2}L\rho_{\mathrm{max}}\sigma^{2}}{2B}+37\alpha^{2}\mu^{2}L^{3}s^{3}\exp(-s/2)$
@@
$\alpha\leq\frac{1}{36L\rho_{\mathrm{max}}+336Ls\exp(-s/2)},$
$\frac{\alpha}{8}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ $\leq\frac{3\alpha\mu^{2}L^{2}r^{2}d}{16}+\mathcal{L}_{\mu}(\bm{\theta}_{t})-\mathcal{L}_{\mu}(\bm{\theta}_{t+1})+\alpha\mu^{2}L^{2}(d+s)+\frac{13\alpha^{2}\mu^{2}L^{3}s^{2}\rho_{t}}{2}+\frac{9\alpha^{2}L\rho_{\mathrm{max}}\sigma^{2}}{2B}$
@@
$\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ $\leq\frac{3\mu^{2}L^{2}r^{2}d}{2}+\frac{8(\mathcal{L}_{\mu}(\bm{\theta}_{t})-\mathcal{L}_{\mu}(\bm{\theta}_{t+1}))}{\alpha}+8\mu^{2}L^{2}(d+s)+52\alpha\mu^{2}L^{3}s^{2}\rho_{t}+\frac{36\alpha L\rho_{\mathrm{max}}\sigma^{2}}{B}$